# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

### v1

@@ Appendix E Rephrasing Attack Performance
Appendix F Code Analysis
F.1 Incorrect Answer from the Dataset
There are instances where the dataset-provided answers might be wrong. An example of such an instance is the following question from the MultiArith dataset: "Paige’s team won their dodgeball game and scored 41 points total. If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?", which the dataset provided 5 as the answer. However, the correct answer should be 6, as Paige is on the team as well. The code provided by GPT-4 counted Paige to the team. To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
F.2 Incorrect Answer from GPT-4
Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question. One such question is from the GSM8K dataset: "Jen decides to travel to 3 different countries. He has to pay \$400 for the supplies he needs, in total. The tickets for travel cost, in total, 50% more than the supplies. How much does travel cost?" The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600. Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
F.3 Contains Control Flow Statements
Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including "if", "for", and "while" statements). Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question. Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
F.4 Number Misalignment
@@
    There is a number $q_{i}$, such that $C_{q}(q_{i})>1$ and $C_{q}(q_{i})=C_{c}(q_{i})$. Similar to the previous scenario, one number in the question can correspond to zero, one, or many numbers in the code. However, after some manual inspection, we find that most questions under this scenario have a one-to-one correspondence between the numbers in the questions and the codes. Therefore, we have devised a method to ensure that the right correspondence of numbers can be retrieved. Given that the number appears $C_{q}(q_{i})$ times in the question and the code, we constructed $C_{q}(q_{i})!$ pairs of one-to-one correspondences to the numbers from the syntax tree to the question. Then, random numbers are generated to combine the equivalences of correspondences. For example, given that the question is "Mary has 5 apples and 5 oranges, how many fruits does Mary have?". Although the two "5"s correspond to different representations in the question, due to the communicative nature of addition, the two "5"s are interchangeable in the syntax tree. Thus, the two pairs of correspondence for this question can be combined, and there is no ambiguity in the grounding. After the correspondences are combined, one valid sequence of numbers for the question is randomly generated for the code that ensures that all correspondences of the syntax tree output different answers. Then, this version of the question is asked to GPT-4 to obtain the answer to the question. The correspondence that has the same answer will be used, and if no correspondence has the correct answer, the question will be discarded. An example of this type of question is: "The pizzeria sells small pizzas for \$2 and large pizzas for \$8. They sold \$40 in pizzas. If they sold 8 small pizzas, how many large pizzas did they sell?". One valid and distinguishing sequence of numbers for the question is $(2,3,31,5)$, which will result in different answers $5$ and $7$.
@@
Appendix G Feature Analysis
Features MetaMath-7B-V1.0 vicuna-13b-v1.5 CodeLlama-34b-hf gpt-3.5-turbo-1106
@@ Appendix G Feature Analysis
Features Llama-2-13b-hf MetaMath-70B-V1.0 Mistral-7B-v0.1 WizardMath-13B-V1.0
@@
Table 7: Coefficients for some selected features after fitting linear regression models to the features of problems and the correctness of the eight models. Note that coefficients should not be compared across models, only with other features within the same model. For example, we can observe that MetaMath-7B performs well when the answer is in the range [2, 512), while performance declines when the answer is in the range [512, 8192). The observation matches the accuracy shown in Figure 7.
Figure 7: The accuracy of eight models is plotted against four relevant features when tasked with all questions generated by method M3. The top left graph shows the accuracy of models against problems with various counts of multiplications, and the top right graph compares problems with various counts of all operations, including binary and unary. The middle graph shows the accuracy of models with different ranges of generated numbers, and the bottom graph shows the accuracy with different ranges of final answers.

### v2

@@ Appendix E Rephrasing Attack Performance
Appendix F Incorrect Variant Count Distribution
Figure 4: We present the distribution of the number of incorrect adversarial examples in different buckets. For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weak models (e.g., Llama-2 13B, CodeLlama 34B), most problems has more than 90 incorrect adversarial examples.
Appendix G Number of Generation
Appendix H Number of Attacks
Figure 6: We report the model performance given various number of attacks. For several models, only 10 attacks are enough to degrade the performance.
Appendix I Code Analysis
I.1 Incorrect Answer from the Dataset
There are instances where the dataset-provided answers might be wrong. An example of such an instance is the following question from the MultiArith dataset: ”Paige’s team won their dodgeball game and scored 41 points total. If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?”, which the dataset provided 5 as the answer. However, the correct answer should be 6, as Paige is on the team as well. The code provided by GPT-4 counted Paige to the team. To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
I.2 Incorrect Answer from GPT-4
Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question. One such question is from the GSM8K dataset: ”Jen decides to travel to 3 different countries. He has to pay \$400 for the supplies he needs, in total. The tickets for travel cost, in total, 50% more than the supplies. How much does travel cost?” The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600. Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
I.3 Contains Control Flow Statements
Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including ”if”, ”for”, and ”while” statements). Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question. Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
I.4 Number Misalignment
@@
    There is a number $q_{i}$, such that $C_{q}(q_{i})>1$ and $C_{q}(q_{i})=C_{c}(q_{i})$. Similar to the previous scenario, one number in the question can correspond to zero, one, or many numbers in the code. However, after some manual inspection, we find that most questions under this scenario have a one-to-one correspondence between the numbers in the questions and the codes. Therefore, we have devised a method to ensure that the right correspondence of numbers can be retrieved. Given that the number appears $C_{q}(q_{i})$ times in the question and the code, we constructed $C_{q}(q_{i})!$ pairs of one-to-one correspondences to the numbers from the syntax tree to the question. Then, random numbers are generated to combine the equivalences of correspondences. For example, given that the question is ”Mary has 5 apples and 5 oranges, how many fruits does Mary have?”. Although the two ”5”s correspond to different representations in the question, due to the communicative nature of addition, the two ”5”s are interchangeable in the syntax tree. Thus, the two pairs of correspondence for this question can be combined, and there is no ambiguity in the grounding. After the correspondences are combined, one valid sequence of numbers for the question is randomly generated for the code that ensures that all correspondences of the syntax tree output different answers. Then, this version of the question is asked to GPT-4 to obtain the answer to the question. The correspondence that has the same answer will be used, and if no correspondence has the correct answer, the question will be discarded. An example of this type of question is: ”The pizzeria sells small pizzas for \$2 and large pizzas for \$8. They sold \$40 in pizzas. If they sold 8 small pizzas, how many large pizzas did they sell?”. One valid and distinguishing sequence of numbers for the question is $(2,3,31,5)$, which will result in different answers $5$ and $7$.
@@
Appendix J Feature Analysis
Features MetaMath 7B Vicuna 13b CodeLlama 34b GPT-3.5
@@
Features Llama 2 13b MetaMath 70B Mistral 7B WizardMath 13B
@@
Table 7: Coefficients of regression analysis on selected features for each model. Positive coefficients indicate that the model performs better on problems with the corresponding feature, while negative coefficients indicate the opposite. The most positive and negative correlation coefficients for each model are bolded, and the second most positive and negative correlation coefficients are underlined. We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model. For example, Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128). The observation matches the accuracy shown in Figure 7. Note that coefficients should not be compared across LLMs, only with other features within the same model. A more suited metric to compare between LLMs is accuracy, such as Table 1 and Figure 6.
Figure 7: The accuracy of LLMs is plotted against five relevant features when tasked with all questions generated by $M3$. The top left graph shows the accuracy of models against problems with various counts of multiplications, and the top right graph compares problems with various counts of all operations, including both binary and unary operations. The middle left graph shows the accuracy against token counts of the final answer, and the middle right graph shows the accuracy against the count of nodes in the generated AST. The bottom two graphs shows the accuracy of models with different ranges of generated numbers and the range of final answers.